OpenStudy (anonymous):

Select the set of equations that represents the following situation: The sum of two numbers is 54, and their difference is nine more than the smaller number. What are the two numbers?

OpenStudy (anonymous):

i got 24=54;54=x-9

OpenStudy (anonymous):

let the larger numbers be x and smaller number be y x+y=54-------1 x-y=y+9-----2 x-y=y+9 subtract y on both sides x-y-y=y-y+9------>x-2y=9 now subtract 2nd equation from 1st x+y=54 -x+2y=-9 ----------- 3y=45 y=45/3 y=15 now plug in y value into one of the equation x+15=54 x=54-15 x=39 so the numbers are 39 and 15